OpenStudy (anonymous):

Temperature Solve for C: F=9/5C+32 please show the steps thank you so much

OpenStudy (jamesj):

This is like many other algebra problems you've seen. You want to 'isolate' C on one side by itself, so what do you think is the first step?

OpenStudy (jamesj):

For example if I asked you to solve this equation for x, what would you do? y = x + 5

OpenStudy (jamesj):

Sorry, you're being too slow in responding. I'll give you the first step, then I have to move on. If F = (9/5)C + 32 then I can subtract 32 from both sides to get F - 32 = (9/5)C + 32 - 32 F - 32 = (9/5)C or (9/5)C = F - 32

OpenStudy (jamesj):

good luck from here